OpenStudy (anonymous):

Add and simplify. log2 4 + log2 16

OpenStudy (anonymous):

log2(4*16)=log2(64)=6

OpenStudy (anonymous):

log 2 2^2+log 2 2^4= 2+4=6

OpenStudy (anonymous):

or think this \(\log_2(4)=x\iff 2^x=4\implies \log_2(4)=2\)

OpenStudy (anonymous):

similarly since \(2^4=16\) you know \(\log_2(16)=4\) then you add \(2+4\)